# THERADIA WoZ: An Ecological Corpus for Appraisal-based Affect Research in Healthcare

### v1

@@ Abstract
We present THERADIA WoZ, an ecological corpus designed for audiovisual research on affect in healthcare. Two groups of senior individuals, consisting of 52 healthy participants and 9 individuals with Mild Cognitive Impairment (MCI), performed Computerised Cognitive Training (CCT) exercises while receiving support from a virtual assistant, tele-operated by a human in the role of a Wizard-of-Oz (WoZ). The audiovisual expressions produced by the participants were fully transcribed, and partially annotated based on dimensions derived from recent models of the appraisal theories, including novelty, intrinsic pleasantness, goal conduciveness, and coping. Additionally, the annotations included 23 affective labels drew from the literature of achievement affects. We present the protocols used for the data collection, transcription, and annotation, along with a detailed analysis of the annotated dimensions and labels. Baseline methods and results for their automatic prediction are also presented. The corpus aims to serve as a valuable resource for researchers in affective computing, and is made available to both industry and academia.
@@ I Introduction
Over the past decade, special attention has been given to the use of Artificial Intelligence (AI) technologies to improve healthcare. One of the rationale behind this craze is that the automation of tasks that do not necessarily require human intervention saves therapists time and mental load, resulting in more efficient practices and less burden for both patients and clinicians. Recent reviews have however evidenced that, even though AI technologies’ performance can rival that of humans in certain assignment tasks, such as serum analysis, or assessment of cardiovascular MRI images, they still struggle to deal with social interaction skills [1, 2].
While the definition of social interaction skills remains a subject of debate [3], they can be summarised as the ability of autonomous agents to maintain equilibrium in a dynamic relationship with another agent, resulting in the development of autonomous relationships [4]. Applied to the context of interactions between AI technology and a patient in healthcare, it corresponds to the capabilities of systems to detect and respond appropriately to signals from patients that may pose a threat to the exchange equilibrium; e.g., patients may express that they feel misunderstood or unsatisfied. It is important to note that patients may convey feelings of being misunderstood through a verbal request, as well as the expression of an affective state. While verbally expressed requests are fairly well understood by today’s AI technologies, it is still difficult for them to accurately detect patients’ affective states from non-verbal signals [5].
@@
Yet using AI technologies for social interactions in healthcare would have far-reaching consequences. A prime illustration of their relevance lies in Computerised Cognitive Training (CCT), which aims to enhance or preserve cognitive functions in patients suffering from cognitive impairment through the repetition of exercises that target specific functions such as attention or memory [7]. Although the effectiveness of CCT has been evidenced by meta-analyses [7, 8, 9, 10], it seems to be conditioned by social interaction during the process [11], which requires the presence of the therapist. Thus, entrusting social interaction to AI technologies would be highly beneficial in making the process autonomous. To move in this direction, it is important to use meaningful data and reliable methods to accurately recognise patients’ affective state in the context of CCT.
In this study, we introduce the THERADIA WoZ corpus, an ecological corpus specifically tailored for the audiovisual detection of affective states in the healthcare domain. The corpus was built to help the development of a virtual assistant supporting CCT sessions at home, for older healthy people and patients [6]. To the best of our knowledge, this is the first audiovisual corpus of affect in healthcare that fully relies on models of appraisal theories, while being accessible to both industrial and academic research communities [^1] The URL will be provided for camera-ready.
@@
A comprehensive analysis of the annotated labels was carried out, enabling the identification of a core set comprising ten affective states in the context of AI technology-human interactions in healthcare. This core set of affective labels was then used along the appraisal dimensions to evaluate the performance of inference models based on audio, visual, and textual data.
The reminder of this paper is organised as follows. We review both the definition and annotation methods of affect according to the appraisal theory, along with the existing corpora in healthcare in section II. The protocol used for the construction of our data set is then presented with the material and the participants in section III. The segmentation and the annotation of the recordings are explained in section IV. A comprehensive analysis of the affective annotations is then reported in section V, and results obtained in their automatic prediction are finally reported in section VI, before concluding in section VII.
@@ II State of the art
In this section, we first review the definition of affective states and their annotation methods through the lens of appraisal theories, and then provide a concise overview of the affect corpora available in healthcare.
@@ II-A Affective states in the lens of the appraisal theory
According to appraisal theories, affective phenomena are defined as multi-component processes aiming to maintain the well-being and survival of individuals in their environment [12, 13]. Roughly speaking, affective phenomena would arise when the meaning of environmental stimuli are – cognitively – appraised as potentially threatening to the equilibrium of individuals in their environment.
@@
Following this appraisal process, the activation of organism’s subsystems, i.e., physiological responses, action tendencies and motor expressions, would be synchronised to help individuals in carrying out the behaviors that contribute to the restoration of their homeostasis. In this perspective, an affective state serves as a marker reflecting how individuals appraise themselves between the plausible deregulation of their homeostasis caused by the environment, and the search for the behavior leading to a return to equilibrium.
The cognitive appraisal process is supposed to be causal in the elicitation of affective episodes [15]. This position assumes a necessary and sufficient relationship would link appraisal to the synchronisation of the organism’s subsystems [16]. In other words, a specific pattern of criteria of the appraisal process would trigger a specific synchronised activation of organism’s subsystems. Conversely, the observation of a specific synchronised activation would allow to predict the pattern of criteria of the appraisal process performed by the experimenter of the affective episode.
Based on these premises, the Tripartite Emotion Expression and Perception model postulates that the way humans perceive affect of others is conditioned by this causal relationship [17]. The idea is that if an individual A is able to determine the affective state of an individual B, this is because A inferred – in part unconsciously – the cognitive appraisal process performed by B on the basis of the apparent motor expressions, e.g., facial expressions, that B has produced. In brief, appraisal theories propose to solve the problem of affective state differentiation based on a $n$-dimensional hyperplane representing the criteria of the appraisal process.
@@ II-B1 Dimensional representations
Therefore, one way of describing individuals’ affective states would be to consider them as points on a plane. Based on this approach, most of the literature have used the Circumplex model to represent them, a Cartesian plane defined by two axes, namely valence – from unpleasant to pleasant – and arousal – from deactivation to activation [18, 19]. However, this approach drew a number of criticisms for its ability to differentiate affective states; fear and anger are both defined by high arousal and negative valence, and the concept of valence itself should be rather seen as a multidimensional representation of several dimensions [20]. It has therefore been stated that more than two dimensions should be taken into account to appropriately distinguish affective states [21]. Thus, appraisal theory represents a sound framework for selecting relevant dimensions in the context of affective state annotation.
@@ II-B2 Categorical representations
Labels are another common way of approximating individual’s affective states. One of the crucial points of such an approach is to select affective labels likely to be relevant in the context of interest. While basic emotion labels, i.e., fear, disgust, happiness, sadness, anger are commonly used to model affective states [22], they are not necessarily the most appropriate in the context of healthcare and AI-patients interactions, as they refer to affect from an evolutionary perspective.
@@ II-B3 Relation between categorical and dimensional representations
Appraisal theories suggest that the use of labels represents a way for humans to characterise the conscious part of the activation states of the organism’s subsystems, i.e., of the cognitive appraisal process, the physiological responses, the action tendencies, and the motor expressions [23, 24], assuming that each affective label would represent a specific pattern of criteria of the appraisal process [17, 25, 26].
@@ II-C Affect corpora for healthcare
Automatic affect analysis can be applied to various aspects of healthcare [27]: to support clinicians for the diagnosis of affective disorders [28], to facilitate the monitoring of patients’ emotions during the therapy [6], which is the focus of this study, or between therapy sessions [29].
Whereas numerous audiovisual corpora are available in the literature for affect recognition in human interactions [30], few of them were acquired in the context of healthcare. There are many reasons for this gap in the literature. One of them concerns the sensitive nature of the data that are protected by law and cannot be easily released. Moreover, populations related to healthcare are often fragile, and setting up experimental protocols to contrast affect corpora can be challenging because of their cognitive fatigue.
Nevertheless, there have been initiatives in the construction of affective corpora supporting clinicians during therapy. For instance, speech interactions between seniors and a virtual assistant playing the role of a health professional coach, were collected for the purpose of the EMPATHIC project [31], which aims to help the elderly maintain their independence as they age [32]. Authors collected $\sim$13k speech segments ($\sim$13h in total) from 136 participants over 64 years and from different countries (Spain, France, and Norway), that were further annotated by three annotators according to five labels: calm, happy, puzzled, tense, and sad, and three dimensions: arousal, valence, and dominance.
@@ III Data collection
The whole data collection protocol of this study has been documented in details, reviewed, and approved by the Ethics Committee for Research in Grenoble Alpes with the reference: CERGA-Avis-2021-1. As the experiment lasted around three hours in average, and participants were asked to consent for the commercial exploitation of their data under the protection of the European General Data Protection Regulation laws, a 20€ gift card was awarded. We also provided reimbursement for travel expanses to participants coming from a certain distance. The protocol used to collect data for the THERADIA WoZ corpus is detailed below.
@@ III-A Participants
A total of 52 healthy older participants (40 females; mean age = 67.9, SD age = 5.1) were recruited based on advertisements published in regional newspapers, i.e., “Le Progrès” and “Le Dauphiné Libéré”. 9 participants diagnosed with MCI (two females; mean age = 75.8, SD age = 2.1) were recruited as patients of clinicians collaborating on the project. Inclusion was based on compliance with the following criteria: fluency in French, normal or corrected-to-normal vision and hearing, and consent prior to the start of the experiment. Regarding the level of education, 40% of healthy older participants had a master’s degree or higher, 33% a bachelor’s degree and 27% a baccalaureate. As for MCI participants, 22% had obtained a master’s degree or higher, 22% a bachelor’s degree, and 56% a National Vocational Qualification (French CAP), or lower.
@@ III-B CCT supported by a Virtual Assistant
All participants carried out a CCT session, which was based on eight exercises selected from the HappyNeuronPro CCT platform [^2] https://www.happyneuron.fr. The exercises involved different cognitive functions such as memory, language, attention, and planning; details on these exercises are provided in section I.A of the supplementary material.
The CCT was supported by a virtual assistant, developed by the Dynamixyz company [^3] Dynamixyz has been acquired by Take-Two Interactive in 2021., and controlled by one co-author of this study acting as a WoZ, cf. Figure 2. More specifically, head movements, gaze, speech, and articulation of the experimenter were captured in real-time to drive the 3D virtual assistant whose rendering was cast to the participant’s screen. The speech was based on a dialogue tree designed to cover a wide range of interactions, envisioned prior to the study. If none of the planned branches made sense within a given dialogue, free intervention was allowed. Eventually, the dialogue tree was updated if the unplanned interaction was deemed relevant. Technical problems that could not be solved by the operator of the virtual assistant were addressed by a second experimenter, who was sitting outside the room for the healthy participants, and inside the room for the MCI participants.
@@ III-C Material
Questionnaires were used to interrogate various psychological, affective and cognitive dimensions of the participants, as well as to evaluate the virtual assistant and the CCT [34]. Experiments were carried out in two rooms of the EMC laboratory of Université Lyon 2, one for the participant, and one for the experimenter.
@@
As pre-tests showed that the video recordings had random freezes that did not show up during the CCT session, we used an iPhone X to provide a stable video recording on the participant’s side, as well as an additional audio recording captured by a far-field omnidirectional microphone. Audio data were sampled at 44.1 kHz with 16 bits, whereas video data were sampled at a constant frame rate of 30 fps with a resolution of 1920x1080 and the YUV420p color format.
@@ III-D Procedure
Concerning the MCI participants, the session began with the completion of the consent form and the above-mentioned questionnaires. Then, after a few instructions, the experimenter left the room so that the CCT session could start. It comprised the realisation of four exercises, each performed twice. The difficulty level of the second exercise was adjusted based on the performance of the first. It was increased in the case of success and decreased in the case of difficulty. Throughout the session, participants were guided and engaged with a virtual assistant under the impression it was operated using AI technology, whereas it was actually operated by a human as a WoZ. The story line of the session started with a welcome dialogue designed to acquaint the participant with the virtual assistant, and to collect some information such as the participant’s mood and motivation. This was followed by the exercise phase divided into five parts: the introductory dialogue detailing each exercise, the exercise completion, the feedback dialogue commenting on participant’s performance, the repetition of the exercise, as well as the feedback dialogue, and the closing dialogue, which ended the session. Once the CCT session was terminated, the whole set-up was revealed.
Concerning the healthy older participants, the procedure was the same, except for the following: First, a session consisted of eight exercises instead of four. Second, healthy older participants were divided into five groups. Twenty of them were assigned to a group without induction. The special feature of this group was that it carried out two CCT sessions, one week apart. This resulted in the set-up being revealed at the end of the second session. The remaining 32 healthy older participants were divided into four induction groups, designed to increase the chances of observing affectively-charged expressions. To this end, induction was used in one out of two of the eight exercises based on the combination of three parameters: Presenting the exercise as easy or hard beforehand; Actually setting the difficulty as easy or hard; And giving feedback from the WoZ as critical or congratulatory. Details on the groups and the inductions are reported in section I.B of the supplementary material.
TABLE I: Overall statistics of the data collected in the THERADIA-Woz.

### v2

@@ Abstract
We present THERADIA WoZ, an ecological corpus designed for audiovisual research on affect in healthcare. Two groups of senior individuals, consisting of 52 healthy participants and 9 individuals with Mild Cognitive Impairment (MCI), performed Computerised Cognitive Training (CCT) exercises while receiving support from a virtual assistant, tele-operated by a human in the role of a Wizard-of-Oz (WoZ). The audiovisual expressions produced by the participants were fully transcribed, and partially annotated based on dimensions derived from recent models of the appraisal theories, including novelty, intrinsic pleasantness, goal conduciveness, and coping. Additionally, the annotations included 23 affective labels drawn from the literature of achievement affects. We present the protocols used for the data collection, transcription, and annotation, along with a detailed analysis of the annotated dimensions and labels. Baseline methods and results for their automatic prediction are also presented. The corpus aims to serve as a valuable resource for researchers in affective computing, and is made available to both industry and academia.
@@ I Introduction
Over the past decade, special attention has been given to the use of Artificial Intelligence (AI) technologies to improve healthcare. One of the rationales behind this craze is that the automation of tasks that do not necessarily require human intervention saves therapists time and mental load, resulting in more efficient practices and less burden for both patients and clinicians. However, recent reviews have evidenced that, even though AI technologies’ performance can rival that of humans in certain specific tasks, such as serum analysis, or assessment of cardiovascular MRI images, they still struggle to deal with social interaction skills [1].
While the definition of social interaction skills remains a subject of debate [2], they can be summarised as the ability of autonomous agents to maintain equilibrium in a dynamic relationship with another agent, resulting in the development of autonomous relationships [3]. In the context of interactions between AI technology and patients in healthcare, it corresponds to the capabilities of systems to detect and respond appropriately to signals from patients that may disrupt the balance of interaction; e.g., patients may express that they feel misunderstood or unsatisfied. While verbally expressed requests are fairly well understood by today’s AI technologies, it is still difficult for them to accurately detect patients’ affective states from non-verbal signals [4].
@@
Yet, the use of AI technologies for social interactions in healthcare would have far-reaching consequences. A prime illustration of their relevance lies in Computerised Cognitive Training (CCT), which aims to enhance or preserve cognitive functions in patients suffering from cognitive impairment through the repetition of exercises that target specific functions such as attention or memory [6]. Although the effectiveness of CCT has been evidenced by meta-analyses [6, 7], it seems to be conditioned by social interaction during the process [8], which requires the presence of the therapist. Thus, entrusting social interaction to AI technologies would be highly beneficial in making the process autonomous. To move in this direction, it is important to use meaningful data and reliable methods to accurately recognise patients’ affective state in the context of CCT.
In this study, we present the THERADIA WoZ corpus, an ecological corpus specifically tailored for the audiovisual detection of affective states in the healthcare domain. The corpus was built to help the development of a virtual assistant supporting CCT sessions at home, for older healthy people and patients [5]. To the best of our knowledge, this is the first audiovisual corpus of affect in healthcare that fully relies on models of appraisal theories, while being accessible to both industrial and academic research communities [^1] The URL will be provided for camera-ready.
@@
A comprehensive analysis of the annotated labels was carried out, enabling the identification of a core set comprising ten affective states in the context of AI technology-human interactions in healthcare. This core set of affective labels was then used along with the appraisal dimensions to evaluate the performance of inference models based on audio, visual, and textual data.
The remainder of this paper is organised as follows. We first review the appraisal theory and annotation methods of affect, along with the existing corpora in healthcare in section II. The protocol used for the construction of our data set is then presented with the material and the participants in section III. The segmentation and the annotation of the recordings are explained in section IV. A comprehensive analysis of the affective annotations is then reported in section V, and results obtained in their automatic prediction are finally reported in section VI, before concluding in section VII.
@@ II State of the art
In this section, we first review the definition of affective states and their annotation methods through the lens of appraisal theories, and then provide a concise overview of corpora of affect available in healthcare.
@@ II-A Affective states in the lens of the appraisal theory
According to appraisal theories, affective phenomena are defined as multi-component processes aiming to maintain the well-being and survival of individuals in their environment [9]. Roughly speaking, affective phenomena arise when the meaning of environmental stimuli are cognitively appraised as potentially threatening to an individual’s equilibrium in their environment.
@@
Following this appraisal process, the activation of the organism’s subsystems, i.e., physiological responses, action tendencies and motor expressions, would be synchronised to help individuals carry out the behaviours that contribute to the restoration of their homeostasis. In this perspective, an affective state serves as a marker reflecting how individuals appraise themselves between the plausible deregulation of their homeostasis caused by the environment, and the search for the behaviour leading to a return to equilibrium.
The cognitive appraisal process is supposed to be causal in the elicitation of affective episodes [11]. This position assumes a necessary and sufficient relationship would link appraisal to the synchronisation of the organism’s subsystems [12]. In other words, a specific pattern of criteria of the appraisal process would trigger a specific synchronised activation of the organism’s subsystems. Conversely, the observation of a specific synchronised activation would allow to predict the pattern of criteria of the appraisal process performed by the experimenter of the affective episode.
Based on these premises, the Tripartite Emotion Expression and Perception model postulates that the way humans perceive the affect of others is conditioned by this causal relationship [13]. The idea is that if an individual $A$ is able to determine the affective state of an individual $B$, this is because $A$ inferred – in part unconsciously – the cognitive appraisal process performed by $B$ on the basis of the apparent motor expressions, e.g., facial expressions, that $B$ has produced. In brief, appraisal theories propose to solve the problem of affective state differentiation based on a $n$-dimensional hyperplane representing the criteria of the appraisal process.
@@ II-B1 Dimensional representations
Therefore, one way of describing individuals’ affective states would be to consider them as points on a plane. Based on this approach, most of the literature have used the Circumplex model to represent them, a Cartesian plane defined by two axes, namely valence – from unpleasant to pleasant – and arousal – from deactivation to activation [14]. However, this approach drew a number of criticisms for its ability to differentiate affective states; fear and anger are both defined by high arousal and negative valence, and the concept of valence itself should be rather seen as a multidimensional representation of several dimensions [15]. It has therefore been stated that more than two dimensions should be taken into account to appropriately distinguish affective states [16]. Thus, appraisal theory represents a sound framework for selecting relevant dimensions in the context of affective state annotation.
@@ II-B2 Categorical representations
Labels are another common way of approximating individual’s affective states. One of the crucial points of such an approach is to select affective labels likely to be relevant in the context of interest. While basic emotion labels, i.e., fear, disgust, happiness, sadness, and anger are commonly used to model affective states [17], they are not necessarily the most appropriate in the context of healthcare and AI-patients interactions, as they refer to affect from an evolutionary perspective.
@@ II-B3 Relation between categorical and dimensional representations
Appraisal theories suggest that the use of labels represents a way for humans to characterise the conscious part of the activation states of the organism’s subsystems, i.e., of the cognitive appraisal process, the physiological responses, the action tendencies, and the motor expressions [18], assuming that each affective label would represent a specific pattern of criteria of the appraisal process [13, 19].
@@ II-C Affect corpora for healthcare
Automatic affect analysis can be applied to various aspects of healthcare [20]: to support clinicians for the diagnosis of affective disorders, to facilitate the monitoring of patients’ emotions during the therapy [5], which is the focus of this study, or between therapy sessions.
Whereas numerous audiovisual corpora are available in the literature for affect recognition in human interactions [21], few of them were acquired in the context of healthcare. There are several reasons for this gap in the literature. One of them concerns the sensitive nature of the data that are protected by law and cannot be easily released. Moreover, populations related to healthcare are often fragile, and setting up experimental protocols to contrast affect corpora can be challenging because of their cognitive fatigue.
Nevertheless, there have been initiatives in the construction of affective corpora supporting clinicians during therapy. For instance, speech interactions between seniors and a virtual assistant playing the role of a health professional coach, were collected for the purpose of the EMPATHIC project [22], which aims to help the elderly maintain their independence as they age. Authors collected $\sim$13k speech segments ($\sim$13h in total) from 136 participants over 64 years and from different countries (Spain, France, and Norway), that were further annotated by three annotators according to five labels: calm, happy, puzzled, tense, and sad, and three dimensions: arousal, valence, and dominance.
@@ III Data collection
The whole data collection protocol of this study has been documented in detail, reviewed, and approved by the Ethics Committee for Research in Grenoble Alpes with the reference: CERGA-Avis-2021-1. As the experiment lasted around three hours on average, and participants were asked to consent for the commercial exploitation of their data under the protection of the European General Data Protection Regulation laws, a 20€ gift card was awarded. Participants could ask for reimbursement of travel expenses, i.e., including public transport fares or petrol costs, for both the journey to and from the experiment site. The protocol used to collect data for the THERADIA WoZ corpus is detailed below.
@@ III-A Participants
A total of 52 healthy senior participants (40 females; mean age = 67.9, SD age = 5.1) were recruited based on advertisements published in regional newspapers such as “Le Progrès” and “Le Dauphiné Libéré”. Additionnaly, 9 participants diagnosed with MCI (two females; mean age = 75.8, SD age = 2.1) were recruited via clinicians collaborating on the project. Inclusion criteria included fluency in French, either normal or corrected-to-normal vision and hearing, and informed consent prior to recordings. Regarding the level of education, 40% of healthy senior participants had a master’s degree or higher, 33% a bachelor’s degree and 27% a baccalaureate. As for MCI participants, 22% had obtained a master’s degree or higher, 22% a bachelor’s degree, and 56% a National Vocational Qualification (French CAP), or lower.
@@ III-B CCT supported by a Virtual Assistant
All participants carried out a CCT session, which was based on eight exercises selected from the HappyNeuronPro CCT platform. The exercises involved different cognitive functions such as memory, language, attention, and planning; details on these exercises are provided in section I.A of the supplementary material.
The CCT was supported by a virtual assistant, developed by the Dynamixyz company, and controlled by one co-author of this study acting as a WoZ, cf. Figure 2. More specifically, head movements, gaze, speech, and articulation of the experimenter were captured in real-time to drive the 3D virtual assistant whose rendering was cast to the participant’s screen. The speech was based on a dialogue tree designed to cover a wide range of interactions, envisioned prior to the study. If none of the planned branches made sense within a given dialogue, free intervention was allowed. Eventually, the dialogue tree was updated if the unplanned interaction was deemed relevant. Technical problems that could not be solved by the operator of the virtual assistant were addressed by a second experimenter, who was sitting outside the room for the healthy participants, and inside the room for the MCI participants.
@@ III-C Material
Questionnaires were used to interrogate various psychological, affective and cognitive dimensions of the participants, as well as to evaluate the virtual assistant and the CCT [24]. Experiments were carried out in two quiet rooms of the EMC laboratory of Université Lyon 2, one for the participant, and one for the experimenter.
@@
As pre-tests revealed random freezes in the video recordings that were not apparent during the CCT session, we used an iPhone X to provide a stable video recording on the participant’s side, along with an additional audio recording captured by a far-field omnidirectional microphone. Audio data were sampled at 44.1 kHz with 16 bits, whereas video data were sampled at a constant frame rate of 30 fps with a resolution of 1920x1080 and the YUV420p color format.
@@ III-D Procedure
Concerning the MCI participants, the session began with the completion of the consent form and the above-mentioned questionnaires. Then, after a few instructions, the experimenter left the room so that the CCT session could start. It comprised the realisation of four exercises, each performed twice. The difficulty level of the second exercise was adjusted based on the performance of the first. It was increased in the case of success and decreased in the case of difficulty. Throughout the session, participants were guided and engaged with the virtual assistant under the impression it was operated using AI technology, whereas it was actually operated by a human as a WoZ. The story line of the session started with a welcome dialogue designed to acquaint the participant with the virtual assistant, and to collect some information such as the participant’s mood and motivation. This was followed by the exercise phase divided into five parts: the introductory dialogue detailing each exercise, the exercise completion, the feedback dialogue commenting on participant’s performance, the repetition of the exercise, as well as the feedback dialogue, and the closing dialogue, which ended the session. Once the CCT session was terminated, the whole set-up was revealed.
Concerning the healthy senior participants, the procedure was the same, except for the following: First, a session consisted of eight exercises instead of four. Second, healthy senior participants were divided into five groups. Twenty of them were assigned to a group without induction. The special feature of this group was that it carried out two CCT sessions, one week apart. This resulted in the set-up being revealed at the end of the second session. The remaining 32 healthy senior participants were divided into four induction groups, designed to increase the chances of observing affectively-charged expressions. To this end, induction was used in one out of two of the eight exercises based on the combination of three parameters: (i) presenting the exercise as easy or hard beforehand; (ii) actually setting the difficulty as easy or hard; and (iii) giving feedback from the WoZ as critical or congratulatory. Details on the groups and the inductions are reported in section I.B of the supplementary material.
TABLE I: Overall statistics of the data collected in the THERADIA-Woz